OpenStudy (anonymous):

Having some trouble with the Limit Definition, See question inside:

OpenStudy (anonymous):

\[y=\sqrt{2x-1}\]

OpenStudy (anonymous):

I know that the limit definition is: \[f(h+x)-f(x)\div h \] But the square root is what is messing me up.

OpenStudy (anonymous):

\[\frac{\sqrt{2(x+h)-1}-\sqrt{2(x)-1}}{h}\]

OpenStudy (anonymous):

oohhh, then multiply by the conjugate, right? I was trying to make \[(2x-1)^{1/2}\] and solve it that way. Thank you.

OpenStudy (anonymous):

you could do that too

OpenStudy (anonymous):

also you can check you work by taking derivative